OpenStudy (anonymous):

Factor completely. 2x^3+2x^2-12x

OpenStudy (anonymous):

first take out 2x =2x(x^2+x-6) =2x(x^2 +3x-2x-6) =2x[x(x+3)-2(x+3)] =2x[(x-2)(x+3)] =2x(x-2)(x+3)